# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v1

@@ 3.3 Route
When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the SDV should follow. In HD map-free settings, however, this is not possible, especially for turning scenarios in intersections where no leading landmark is available. In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues. The lane center lines constructed by landmarks are mainly used in cruise and lane change scenarios, while in turning scenarios, the starting and ending positions are recorded in advance, once vehicles approach the intersection, the human commuting trajectories will serve as route. More details can be found in the appendix.
@@ 3.4 Prediction
The decoder, as mentioned in 3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box. This grid is trained with pixel-wise focal loss inspired from [15]:
@@
In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid. For the agent-level prediction, we follow the interactive design. Since the output trajectories are mainly adopted in sampling for SDV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego. In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior. Those landmarks indicate areas where interaction is most likely to happen. The agents’ trajectories will be decoded conditioned on that lane prior:
@@ 3.5 Trajectory Evaluation
The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories. We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost maps for scoring candidate trajectories. We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map. Our evaluator decodes a space-time cost map $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster. The cost of a trajectory is defined as the sum of costs at its space-time way-points. The best candidate trajectory is the one with the lowest cost according to the learned cost:
@@ 3.6 Trajectory Generation
The sampling and generating algorithms can be mainly categorized into reference line-based and reference line-free. Methods incorporating reference information can generate long-term consistent outputs in the loops. Since this information rarely changes over time, bridging consecutive frames promotes cross-frame stability. However, scenarios still need to be handled where reference-related information can hardly be obtained. To enhance the scalability of our system, we consider trajectory from the following methods. In our implementation, curve and retrieval-based samplers are reference line-free, while the others rely on reference information constructed by visual landmarks or commuting trajectories.
Curve Based Sampler
As the most basic sampler, it has been adopted by a few previous methods[48]. We follow [48] with little modifications for comfort concerns. We take into account the initial states of a Self-Driving Vehicle(SDV), such as velocity, acceleration, heading, and road curvature where the SDV is located, and then sample a combination of straight lines, circles, and clothoid curves as trajectories with the given initial states and random end states within a pre-defined range. The curve-based sampler serves as the baseline sampling algorithm in our experiment. The details can be found in the appendix.
Retrieval Based Sampler
Following prior work that uses data-driven trajectory sampling techniques in mapless driving scenarios [3], we create an expert trajectory dataset by clustering manual driving data. Once this dataset is constructed during pre-processing, SDV can retrieve trajectories with high similarity according to its current motion states. The construction and retrieval details can be found in the appendix.
@@
Qi added The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the SDV. When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process. The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in longitude. In our paradigm, they are both available as depicted by Sec. 3.4 Sec. 3.1. In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39].Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude. We illustrate the detail in the appendix.
@@
$f_{plan}=[f_{scene}||f_{route}||f_{SDV}]$ (7)
@@
The above imitation generator is trained separately from our evaluator by different headers. With the attempt to generate trajectory with minimal cost, we propose a GAN base planner with modification to the training paradigm that bridges the isolated optimization targets, so that Eq. 4 can be solved by learning. For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize below function while the discriminator tries to maximize it.
@@
$log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly. In our paradigm, we replace the discriminator with an evaluator. But different from the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator output score to measure the discrepancy between candidates quantitatively. So the optimization target for evaluator Eq. 5 still holds, but the learning target of the imitation planner becomes:
@@
Here m is an hidden variable represent the manual in planning.
@@ 4 Experiments
We construct different baseline variants of our models in both open-loop and closed-loop for ablative studies and compare them with state-of-the-art alternative methods quantitatively. We gather in-house datasets for training and open-loop settings. It contains a wide range of driving scenarios in a densely populated urban environment. We collected the record files in 15 hours, which captured the perception output, the SDV trajectory, and route info. All the records are post-processed to obtain 200k clips in total. By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future. We partition the dataset into 16w for training and 4w for testing. To validate the potential of our method, we disable the functionality in the safety layer as well as the optimizer and safety backup subsystem in other methods. More details on protocols are provided in the appendix.
@@ 4.1 Prediction
We focus on the $L2$ distance and hit rate for the open-loop prediction. Preview SOTAs are evaluated in our HD map-free setting.
@@ 4.2 Open-loop Planning
For the open-loop planning, we focus on two evaluation metrics: $L2$ error and collision rate, and align the planning horizon to 3.0s for a fair comparison. We compute the $L2$ error between the planned trajectory and the human driving trajectory, and evaluate how often the SDV would collide with other agents on the road. Detailed comparison with previous methods is shown in Tab. Note that the Vanilla algorithm is penalized based on how much the trajectory deviates from the ground truth, thus it achieves the lowest L2 error but largest collision rates in all prediction horizons.
@@ 4.3 Closed-loop Experimental
We first evaluate the necessity of different sampler on closed-loop simulation by ablation study. Then, find the best combination to evaluate in the real-world. The results of ablation study illustrated appendix.
We extensively tested andn appendix. compare the above methods in the real world, in densely populated urban area, under the supervision of human safety drivers. The driving route used for our experiments is shown on appendix. During the 100+ mile public road testing, the model successfully performed a wide variety of challenging maneuvers including lane-following, merging, yielding to pedestrians or nudging around parked cars.
@@ Appendix A Appendix
A.1 Scene Rasterization
@@
The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections. Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect. To distinct SDV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history. The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness. we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route. Since we explore route from different source, we cast different routing form into one single uniformed *** All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
A.2 Instance Vectorization
A.3 Model Structure
A.4 Route
@@
As indicated previously, the route plays a critical role in guiding SDVs along roadways. Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
@@
The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the SDV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
@@
However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes. Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions. Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the SDV to change lanes.
@@
Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the SDV lacking appropriate instructions to navigate intersections. To address this, the starting and ending positions for turning are recorded. As the SDV approaches intersections, human-driving trajectories are employed as the route to guide the SDV through the intersections.
A.5 Curve Based Sampler
@@
In this appendix, we present the sampling procedure with more details. Vehicle trajectories can be categorized into two primary types: straight and turning. And turning trajectories encompass lane changes and turning at intersections. Lane-changing trajectories are characterized by variable curvature over time, while turning at intersections follows circular-like curves. As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.

### v2

@@ 3.3 Route
When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the ADV should follow. However, in HD map-free settings, this is not possible, especially for turning scenarios in intersections where no leading landmark is available. In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues. The lane center lines constructed by landmarks are mainly used in cruise and lane change scenarios, while in turning scenarios, the starting and ending positions are recorded in advance, once vehicles approach the intersection, the human commuting trajectories will serve as route. More details can be found in the appendix.
@@ 3.4 Prediction
The decoder, as mentioned in Sec. 3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box. This grid is trained with pixel-wise focal loss inspired from [15]:
@@
In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid. For the agent-level prediction, we follow the interactive design. Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego. In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior. Those landmarks indicate areas where interaction is most likely to happen. The agents’ trajectories will be decoded conditioned on that lane prior:
@@ 3.5 Trajectory Evaluation
The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories. We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost volumes for scoring candidate trajectories. We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map. Our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster. The cost of a trajectory is defined as the sum of costs at its space-time way-points. The best candidate trajectory is the one with the lowest cost according to the learned cost:
@@ 3.6 Trajectory Generation
The sampling and generating algorithms can be mainly categorized into reference line-based and reference line-free. Methods incorporating reference information can generate long-term consistent outputs in the loops. Since this information rarely changes over time, bridging consecutive frames promotes cross-frame stability. However, scenarios still need to be handled where reference-related information can hardly be obtained. To enhance the scalability of our system, we follow [48, 3] to build curve and retrieval-based sampler in reference-line free and illustrate them in the appendix. The proposed lattice sampler, imitation planner, and GAN-based planner will be explained in this section. They incorperate reference info constructed by visual landmarks or commuting trajectories.
@@
The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the ADV. When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process. The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in longitude. In our paradigm, they are both available as depicted by Sec. 3.4 Sec. 3.1. In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39]. Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude. We illustrate the detail in the appendix.
@@
$f_{plan}=[f_{scene}||f_{route}||f_{ADV}]$ (7)
@@
The above imitation generator is trained separately from our evaluator by different headers. With the attempt to generate trajectory with minimal cost, we propose a GAN base planner with modification to the training paradigm that bridges the isolated optimization targets, so that Eq. 4 can be solved by learning. For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize the following function while the discriminator tries to maximize it.
@@
$log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly. In our paradigm, we replace the discriminator with an evaluator. But unlike the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator outputs a score to measure the discrepancy between candidates quantitatively. So the optimization target for evaluator Eq. 5 still holds, but the learning target of the imitation planner becomes:
@@
Here, m is a hidden variable representing the manual in planning.
However, the trajectory evaluation Eq. 14 involve index trajectory points in the learned cost volume $C_{xy}^{t}=dec(f_{ras})$, which is non-differentiable, thus cannot be back propagated to the generator. To address this issue, we use a soft function to get a meaningful gradient. The grid_sample operator[1] is adopted here to obtain the gradients via interpolation between values in our cost volume.
@@ 4 Experiments
We develop various baseline variants of our models under both open-loop and closed-loop configurations for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies. We gather in-house datasets for both training and open-loop evaluations, encompassing an extensive array of driving scenarios within densely populated urban settings. To substantiate the feasibility of the data-driven planners, we deactivate the functionalities in the safety layer as well as the optimizer and safety backup subsystems in comparative methodologies such as optimal control in UniAD[22], etc.
@@ 4.1 Prediction
For the prediction task, prior state-of-the-art methodologies are appraised within HD map-free context, wherein solely the observable landmarks are furnished to the predictors. We measure the performance of methods on root mean squared error and hit-rate overall future timesteps $t\in{1,2,3}$ seconds. In Tab. 1, we compare all methods using the same inputs. In general, our multi-modal design with interactive prior outperformed our reimplemented TNT[50] with 16 candidates, while the uni-modal version shared similar results with the current state-of-the-art.
Method Mean L2 / hit-rate $<$ 1m Avg.
1s 2s 3s
VectorNet[13] 0.75/0.80 1.72/0.40 2.85/0.22 1.34
GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35
TNT[50] 0.65/0.78 1.01/0.57 1.72/0.34 0.90
GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
Table 1: Prediction Results in uni-modal and multi-modal.
Method trajectory sources L2(m) Collision(%)
curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s
UniAD[22] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78
VAD[24] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69
NMP[48] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05
MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97
P3[38] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04
GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52
GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48
GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07
GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
Table 2: Open loop planning results of imitation based method, sampling based method and GAD.
Method comfort event in 30Km efficiency MPI
Lat. accel. Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute
UniAD[22] 0.351 3.12 3.68 - - - - 15.20 0.31
ST-P3[21] 0.857 4.54 8.53 4 3 14 3 8.62 1.36
GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72
GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75
GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75
GAD-GAN 0.511 1.15 3.52 3 1 2 1 5.51 4.28
DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
Table 3: Road Test Outcomes: Data were systematically collected during a 30 km road test, encompassing the computation of average time per loop. Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded. The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
@@ 4.2 Open-loop Planning
For open-loop planning, our analysis prioritizes two pivotal evaluation metrics: the $L2$ error and the collision rate. The planning horizon is standardized at 3.0 seconds to ensure equitable comparison. We ascertain the $L2$ error by quantifying the deviation of the planned trajectory from the human driving trajectory and assess the frequency of potential collisions involving the ADV with other road agents. A comprehensive comparison with antecedent methodologies is delineated in Tab. 2. It is noteworthy that the imitation strategy incurs penalties proportional to the trajectory’s divergence from the empirical ground truth, resulting in minimal $L2$ errors across all examined horizons.
@@ 4.3 Closed-loop Experimental
Initially, we conducted an ablation study to assess the necessity of various samplers within a closed road environment. Subsequently, we identified the optimal sampler combination for further evaluation on open urban roadways. The outcomes of this ablation study are delineated in the appendix. Comprehensive real-world testing of these methodologies was conducted in densely populated urban settings, under the vigilant supervision of human safety drivers. The specific driving route utilized for our experiments is depicted in the appendix. Throughout the extensive public road testing, spanning over 30 kilometers, the model adeptly executed a diverse array of complex maneuvers, such as lane-following, merging, yielding to pedestrians. The results of these road tests are documented in Tab. 3
It is noteworthy that methodologies exhibiting superior imitation metrics do not invariably demonstrate enhanced performance in closed-loop evaluations. This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks. Nonetheless, the multi-modal imitation generator and GAN-based generator significantly enhance the system’s performance in both open-loop and closed-loop settings, primarily in terms of the planner’s intelligence and the continuity between consecutive frames. The efficiency and comfort metrics both show improvements with the integration of data-driven generators. Closed-loop results suggest that multimodality mitigates the brittleness inherent in imitation learning, while the lattice sampler further enhances the robustness of the entire system.
To rigorously evaluate the efficacy of the proposed methodology as a substitute for off-the-shelf planners, we conducted a comparative analysis between the GAD and a conventional modularized, user-level, rule-based system, referred to as DPQP in Tab. 3. Notwithstanding, a discernible performance discrepancy persists in terms of system’s robustness and comfortableness. We posit that increase training data size could ameliorate this disparity.
@@ 5 Conclusions
We introduce a sophisticated DL pipeline from prediction to motion planning in HD map-free setting, augmented with an enhanced safety assurance. Our investigations have yielded substantial enhancements over prevailing data-driven methodologies within the complex urban environments of Shanghai. This paradigm enables the safety deployment of learned planners in real-world scenarios, offering interpretability and leveraging their capacity to evolve with accumulating data, thereby managing scenarios of greater complexity than those addressed by traditional rule-based systems. Future work may integrate generative learning techniques such as Generative Adversarial Networks (GAN), Conditional Variational Autoencoders (CVAE), and diffusion models to further augment the data-driven capabilities in motion planning. Furthermore, evidence suggests that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners. It is recommended that subsequent research in motion planning incorporate the imitation metric to gauge the intelligence domain of planners and to develop additional metrics for closed-loop evaluation.
@@ Appendix A Appendix
A.1 Perception Outputs
At each timestamp, measured states for each tracked entity are provided in terms of positions, velocities, and accelerations. Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead. Stop lines are also essential when the ADV approaches the junction area. The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form. The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status, characterized as non-parametric.
A.2 Scene Rasterization
@@
The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections. To capture the past motion of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers. Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect. To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history. The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness. we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route. All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
A.3 Instance Vectorization
Parametric information such as surrounding vehicle positions, lane center lines and lane landmarks, are derived with the form of discretized points to depict the traffic scenes. Additionally, route information is obtained to facilitate the ADV in navigating to its intended destination. Given that the vehicle’s historical positions are temporally continuous, whereas the lane center lines, lane landmarks, and route information are spatially continuous, they are structured as vectors $\mathbf{v}$ that connect adjacent points. Specifically, the $i$th vector can be defined as
$\mathbf{v}_{i}=[p_{i},p_{i+1},a_{i}]$ (17)
where $p_{i}$ represents the $i$th point, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
Similar to the process of scene rasterization, traffic lights are considered as static agents and can be vectorized using unchanged vectors. Each vector can be seen as a node in a graph. Subsequently, a sequence of nodes belonging to the same category will be combined to form an instance, which will then be interconnected within the graph.
A.4 Model structure
Concurrently, the vectorization branch processes parametric instance information through a PointNet subgraph, which includes three edge convolutions. The output from this branch is then passed through a downstream Transformer encoder to obtain the instance encoding. Note that the rasterization branch has more layers than the vectorization branch as the input is much denser than the vectorized graph.
Following the extraction of the scene embedding, features by oriented crop around the agents are obtained. These features are then processed through a sequence of max-pooling and a multi-layer perceptron (MLP). The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
A.5 Route
@@
As indicated previously, the route plays a critical role in guiding ADVs along roadways. Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
@@
The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
@@
However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes. Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions. Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the ADV to change lanes.
@@
Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections. To address this, the starting and ending positions for turning are recorded. As the ADV approaches intersections, human-driving trajectories are employed as the route to guide the ADV through the intersections.
A.6 Curve Based Sampler
@@
In this appendix, we present the sampling procedure with more details. We follow [48] with little modifications for comfort concerns. Vehicle trajectories can be categorized into two primary types: straight and turning. And turning trajectories encompass lane changes and turning at intersections. Lane-changing trajectories are characterized by variable curvature over time, while turning at intersections follows circular-like curves. As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.